Find (g of f)(x) and domain of (g of f)
f(x)=4x+5
g(x)= 2/x
(g of f)(x)= Simplified answer...
Domain of (g of f)...

Answer by stanbon(75887)

Find (g of f)(x) and domain of (g of f)
f(x)=4x+5
g(x)= 2/x
--------------------
g[f(x)] = g[4x+5] = 2/(4x+5)
---------------
The denominator cannot be zero.
Domain: All Real Numbers except x = -5/4
==============================================
Cheers,
Stan H.
==============